OpenStudy (anonymous):

Translate to a system of equations and solve. A cyclist rode 3 hours with a tail wind of 6 km/h. The return trip against the head wind took 5 hours. Find the cyclists speed in the still air

OpenStudy (anonymous):

d=r*t, where r is the speed in still air the distance there and back is the same, so (r + 6)*3 = (r - 6)*5 solve for r

OpenStudy (anonymous):

@ashleyodom your question is the same, just replace the times with those for your problem and the 6 with the 40...make sure the + 40 is with the tail-wind time and the - 40 is with the head-wind time.

OpenStudy (anonymous):

oh okay thanks ! wait do you mind writing the equation down like how you did for this one .. thx @mtbender74

OpenStudy (anonymous):

thanks ! @mtbender74

OpenStudy (anonymous):

(r + 40)*3 = (r - 40)*4